# I-MCTS: Enhancing Agentic AutoML via Introspective Monte Carlo Tree Search

### v3

@@ Abstract
Recent advancements in large language models (LLMs) have shown remarkable potential in automating machine learning tasks. However, existing LLM-based agents often struggle with low-diversity and suboptimal code generation. While recent work Chi et al. (2024) has introduced Monte Carlo Tree Search (MCTS) to address these issues, limitations persist in the quality and diversity of thoughts generated, as well as in the scalar value feedback mechanisms used for node selection. In this study, we introduce Introspective Monte Carlo Tree Search (I-MCTS), a novel approach that iteratively expands tree nodes through an introspective process that meticulously analyzes solutions and results from parent and sibling nodes. This facilitates a continuous refinement of the node in the search tree. Furthermore, we integrate a Large Language Model (LLM)-based value model to facilitate direct evaluation of each node’s solution prior to conducting comprehensive computational rollouts. A hybrid rewarding mechanism is implemented to seamlessly transition the Q-value from LLM-estimated scores to actual performance scores. This allows higher-quality nodes to be traversed earlier. Applied to the various ML tasks, our approach demonstrates a 6% absolute improvement in performance compared to the strong open-source AutoML agents, showcasing its effectiveness in enhancing agentic AutoML systems [^1] Resource available at https://github.com/jokieleung/I-MCTS.
@@ 1 Introduction
Recent advances in large language models (LLMs) have opened new frontiers in automating machine learning (AutoML) tasks Huang et al. (2024); Chan et al. (2024). Compared to the traditional AutoML frameworks Feurer et al. (2015); Tang et al. (2024), the LLM-based agents Sun et al. (2024) emerge as promising direction due to their remarkable capabilities on code generation Jimenez et al. (2023), neural architecture design Zheng et al. (2023) and model training Chi et al. (2024). Overall, these LLM agent-based AutoML systems Schmidt et al. (2024); Hong et al. (2024a); Li et al. (2024); Chi et al. (2024) typically input with a natural language description on the dataset and the problem, after which the system directly generates a solution in an end-to-end manner. While recent works by Schmidt et al. (2024) and Hong et al. (2024a) have made significant strides in automating the machine learning workflow, replicating the adaptive and strategic behavior of expert human engineers remains a significant challenge. The primary limitation lies in their search processes, which typically involve only a single pass or trial. This constraint significantly hinders the generation of diverse and highly optimized workflows, a capability that human experts excel at through iterative refinement and strategic decision-making. Recent work by  Chi et al. (2024) introduced Tree-Search Enhanced LLM Agents (SELA), leveraging Monte Carlo Tree Search (MCTS) to expand the search space. However, the search space is constrained by its reliance on a pre-generated, fixed insight pool derived from the initial problem description and dataset information. This static nature inherently limits the diversity and adaptability of the search tree. Moreover, SELA encounters significant limitations in improving the overall quality of solutions when relying exclusively on scalar experimental performance feedback. This reliance on simplistic scalar feedback mechanisms impedes the efficient identification and navigation of optimal pathways in complex machine learning tasks, where flexible re-assessment and adaptive strategies are usually needed.
@@
  • Extensive experiments across diverse ML tasks demonstrate our method’s superiority, achieving 6% absolute performance gains over state-of-the-art baselines while maintaining computational efficiency.
@@ 2.3 Hybrid Reward Mechanism
The primary objective of the evaluation phase is to assess the reward for the current node. For machine learning tasks, a conventional approach Schmidt et al. (2024); Chi et al. (2024) involves utilizing the performance metric on the development set as the reward. To derive a performance score of the given node involves rolling out from an intermediate state to a terminal state, which is quite costly since processes such as model training cost significant computational time. This limits the efficiency of node exploration. To address these challenges, we introduce a hybrid reward mechanism for node evaluation which combines LLM-estimated evaluations with actual performance scores to guide the search process more efficiently. For each node $x$, we employ an LLM-based value model $M_{\text{value}}$ to predict its potential performance. The value model takes as input the solution $\sigma_{\text{sol}}(x)$ and outputs an estimated score $s_{\text{LLM}}(x)$. This score is based on a comprehensive set of machine learning evaluation criteria, which is detailed in Appendix H.2. The LLM-estimated evaluation allows us to assess the quality of a node’s solution before conducting computationally expensive rollouts. This early evaluation helps prioritize nodes with higher potential, improving the efficiency of the search process. After the "simulation" of a node $x$, we obtain the actual performance score $s_{\text{actual}}(x)$ based on the development set, which reflects the true performance of the node’s solution. To smoothly transition from LLM-estimated values to actual performance scores, we implement an adaptive reward blending strategy. The Q-value $Q(x)$ for a node $x$ is updated as follows:
@@ 2.4 Tree Search Process
Upon simulation completion, the Q value is propagated backwards through the tree structure. This process updates the value and visit count of each parent node, from the simulated node to the root. Consequently, nodes associated with superior solutions receive higher priority in subsequent iterations. This allows nodes representing more promising solutions to be prioritized in future rollouts. Similar to Chi et al. (2024), I-MCTS implements a state-saving mechanism that caches the stage code for each node, which allows it to reuse previously generated code when a new configuration shares components with existing ones.
@@ 3 Experiment
For fair comparison, we follow Chi et al. (2024) to benchmark our method on 20 machine learning datasets, including 13 classification tasks and 7 regression tasks from the AutoML Benchmark (AMLB)  Gijsbers et al. (2024) and Kaggle Competitions. Five baselines are included for comparison, including AutoGluon Tang et al. (2024), AutoSklearn Feurer et al. (2022), AIDE Schmidt et al. (2024), Data Interpreter Hong et al. (2024a), SELA Chi et al. (2024). We directly used the results reported in Chi et al. (2024). Experiments are conducted based on the MetaGPT Hong et al. (2024b) framework [^2] https://github.com/geekan/MetaGPT. More experimental details please refer to the Appendix D.
@@ 3.3 Case Study
We conduct a comprehensive visualization analysis of the tree search process using the kick dataset. As show in Appendix I, our empirical results reveal that I-MCTS effectively leverages the introspective information from previous nodes to generate task-specific and actionable insights, while SELA exhibit significant homogeneity and lack specificity. Furthermore, our hybrid reward mechanism enhances the exploration efficiency, enabling more effective identification of high-quality nodes within the same computational budget.
@@ 4 Conclusion
While the proposed I-MCTS approach demonstrates significant improvements in AutoML agent performance, several limitations remain that warrant further investigation. First, the computational overhead of the introspective process, although beneficial for enhancing decision-making, introduces additional resource requirements. This limits the scalability of the method, particularly in scenarios where computational resources are constrained. Future work should explore optimizations to reduce this overhead and investigate how the approach scales with increased computational power.
Second, the current evaluation of I-MCTS is primarily focused on structured data and tabular ML tasks. Its effectiveness on more complex and heterogeneous data types, such as image and text data, remains unexplored. Extending the application of I-MCTS to these domains could provide valuable insights into its generalizability and robustness across diverse machine learning tasks, such as Vison, NLP and RL tasks.
@@ Appendix A Algorithm
0:  Problem description $p$, dataset $D$, metadata $d$, LLM $LLM$, rollout number $k$, blending constant $\gamma$.
@@
12:  $x_{\text{best}}\leftarrow\underset{x\in\text{Tree}}{\text{argmax}}(Q(x))$
12:  $\sigma_{\text{sol}}(x_{\text{best}})$
@@ Appendix B Detailed Results
Click_prediction_small 26.6 26.6 40.2 40.3 26.1 39.4 12.9 13.9 23.2 27.4 29.1 30.2
GesturePhaseSegmentationProcessed 69.3 69.3 67.2 68.4 56.3 68.1 60.1 64.4 67.9 69.2 67.2 68.5
Moneyball 24.3 24.3 13.1 13.8 23.8 24.6 9.5 24.5 21.9 24.5 41.6 41.7
SAT11-HAND-runtime-regression 12.6 12.6 10.3 10.3 12.0 12.1 11.4 11.9 12.2 12.5 21.7 24.6
boston 39.8 39.8 19.5 19.6 40.5 41.3 37.0 38.6 40.1 41.4 62.5 62.7
colleges 88.3 88.3 2.1 2.1 86.0 87.8 87.5 87.7 87.8 87.8 94.1 94.3
concrete-strength 28.3 28.3 17.4 17.9 28.3 28.3 28.8 29.6 28.2 28.2 47.3 48.2
credit-g 50.5 50.5 35.1 44.0 21.6 48.4 48.1 53.2 50.9 52.7 50.9 51.1
diamonds 13.8 13.8 8.7 8.7 13.7 13.7 13.5 13.6 13.7 13.8 26.8 26.8
house-prices 9.0 9.0 2.0 2.0 8.9 8.9 8.5 9.0 8.9 9.0 15.8 18.4
icr 76.2 76.2 70.4 79.2 31.7 35.9 57.8 60.6 78.7 79.2 72.3 79.0
jasmine 84.3 84.3 84.4 84.7 83.6 84.6 77.8 83.5 85.4 86.2 84.2 84.3
kc1 38.3 38.3 43.5 45.0 40.8 42.6 38.1 41.2 42.2 43.1 44.8 45.6
kick 39.6 39.6 41.8 42.1 14.9 38.6 2.8 4.2 35.9 38.7 36.2 42.4
mfeat-factors 96.7 96.7 97.1 97.5 94.4 94.5 93.0 96.0 95.7 96.2 94.4 95.4
segment 93.5 93.5 92.7 93.1 91.7 92.2 91.7 92.6 93.8 94.4 91.1 93.7
smoker-status 78.0 78.0 78.6 78.9 74.8 76.3 77.3 81.5 82.4 91.5 79.9 91.5
software-defects 51.5 51.5 61.1 61.7 49.7 49.8 54.5 57.3 52.2 53.3 60.5 60.7
titanic 78.9 78.9 76.2 78.9 81.2 83.7 76.0 78.5 78.8 79.7 76.1 78.5
wine-quality-white 65.4 65.4 60.7 61.4 62.9 65.1 61.2 61.6 65.3 66.0 63.3 65.5
Overall NS % $\uparrow$ 53.2 53.2 46.1 47.5 45.5 51.8 47.4 50.2 53.3 54.7 58.6 59.8
@@
Appendix C Evaluation Metrics
Appendix D Experimental Details
All datasets are split into training, validation, and test sets with a 6:2:2 ratio. For the LLM-based Agents (I-MCTS, SELA, Data Interpreter, AIDE, etc), we employ a consistent initial task prompt across all methods  [^3] Details can be refer to our code repository https://anonymous.4open.science/r/I-MCTS. This prompt encompasses the dataset name, target column, and evaluation metric. We choose Qwen2.5-72B-Instruct as our foundation LLM, we set the temperature parameter to 0.5. $\alpha_{\text{explore}}$ is set to 2. AIDE conducts 10 iterations per execution, while I-MCTS, SELA perform 10 rollouts. For I-MCTS and SELA, we employ Data Interpreter as the experimenter. The value of $\gamma$ in HRM is set to 0.2 across all datasets. This value is to control the decay rate of the LLM estimated score. Larger $\gamma$ makes the influence of the LLM estimated to persist longer, while smaller $\gamma$ allows the actual performance results to dominate the hybrid reward score more quickly. $\alpha_{explore}$ is set to 2. This aims to balance exploration and exploitation in the method’s search strategy. Following Chi et al. (2024), each method, except for AutoGluon, is run three times for each dataset. AutoGluon, being deterministic, is run only once with its default settings. AutoSklearn is also run with default settings, limited to 600 seconds per task.
Appendix E Datasets
@@
Appendix F Cost-efficiency analysis
Appendix G Related work
Appendix H Prompts
H.1 Introspection Node Expansion Prompt
@@
You are an AI assistant tasked with analyzing a machine learning solution and proposing new insights to IMPORVE its performance. Given the current plan, solution code, Evaluation Metrics and performance on development score, suggest innovative approaches to ENHANCE and IMPORVE the performance.
@@
Based on this information, please propose ONE new insight w.r.t. the following specific stage of the machine learning pipeline to IMPORVE the overall performance. Your insight should be specific, actionable, and have the potential to IMPORVE the model’s performance.
@@
"insight": "ONE insight that specific, actionable, and have the potential to IMPORVE the model’s performance."
H.2 Node Evaluation Prompt
@@
Appendix I Case Study
@@
I.1 Overview of I-MCTS’s search process
@@ 99 Simulated: False
I.2 Overview of baseline SELA’s search process

### v4

@@ Abstract
Recent advancements in large language models (LLMs) have shown remarkable potential in automating machine learning tasks. However, existing LLM-based agents often struggle with low-diversity and suboptimal code generation. While recent work Chi et al. (2024) has introduced Monte Carlo Tree Search (MCTS) to address these issues, limitations persist in the quality and diversity of thoughts generated, as well as in the scalar value feedback mechanisms used for node selection. In this study, we introduce Introspective Monte Carlo Tree Search (I-MCTS), a novel approach that iteratively expands tree nodes through an introspective process that meticulously analyzes solutions and results from parent and sibling nodes. This facilitates a continuous refinement of the node in the search tree. Furthermore, we integrate a Large Language Model (LLM)-based value model to facilitate direct evaluation of each node’s solution prior to conducting comprehensive computational rollouts. A hybrid rewarding mechanism is implemented to seamlessly transition the Q-value from LLM-estimated scores to actual performance scores. This allows higher-quality nodes to be traversed earlier. Applied to the various ML tasks, our approach demonstrates a 4% absolute improvement in performance compared to the strong open-source AutoML agents, showcasing its effectiveness in enhancing agentic AutoML systems [^1] Resource available at https://github.com/jokieleung/I-MCTS.
@@ 1 Introduction
Recent advances in large language models (LLMs) have opened new frontiers in automating machine learning (AutoML) tasks Huang et al. (2024); Chan et al. (2024). Compared to the traditional AutoML frameworks Feurer et al. (2015); Tang et al. (2024), the LLM-based agents Sun et al. (2024) emerge as promising direction due to their remarkable capabilities on code generation Jimenez et al. (2023), neural architecture design Zheng et al. (2023) and model training Chi et al. (2024). Overall, these LLM agent-based AutoML systems Schmidt et al. (2024); Hong et al. (2024a); Li et al. (2024); Chi et al. (2024) typically input with a natural language description on the dataset and the problem, after which the system directly generates a solution in an end-to-end manner. While recent works by Schmidt et al. (2024) and Hong et al. (2024a) have made significant strides in automating the machine learning workflow, replicating the adaptive and strategic behavior of expert human engineers remains a significant challenge. The primary limitation lies in their search processes, which typically involve only a single pass or trial. This constraint significantly hinders the generation of diverse and highly optimized workflows, a capability that human experts excel at through iterative refinement and strategic decision-making. Recent work by  Chi et al. (2024) introduced Tree-Search Enhanced LLM Agents (SELA), leveraging Monte Carlo Tree Search (MCTS) to expand the search space. However, the search space is constrained by its reliance on a pre-generated and fixed insight pool derived from the initial problem description and dataset information. This static nature inherently limits the diversity and adaptability of the search tree. Moreover, SELA encounters significant limitations in improving the overall quality of solutions when relying exclusively on scalar experimental performance feedback. This reliance on simplistic scalar feedback mechanisms impedes the efficient identification and navigation of optimal pathways in complex machine learning tasks, where flexible re-assessment and adaptive strategies are usually needed.
@@
  • Extensive experiments across diverse ML tasks demonstrate our method’s superiority, achieving 4% absolute performance gains over state-of-the-art baselines while maintaining computational efficiency.
2 Related work
At the intersection of multi-agent systems, Fleet of Agents Arora et al. (2024) employs genetic particle filtering to coordinate LLM agents, outperforming Tree-of-Thoughts Yao et al. (2023) methods in efficiency and accuracy for puzzle-solving tasks. Qu et al. (2024) extended these concepts through recursive introspection, enabling agents to systematically self-improve via structured reflection cycles.
@@ 3.3 Hybrid Reward Mechanism
The primary objective of the evaluation phase is to assess the reward for the current node. For machine learning tasks, a conventional approach Schmidt et al. (2024); Chi et al. (2024) involves utilizing the performance metric on the development set as the reward. To derive a performance score of the given node involves rolling out from an intermediate state to a terminal state, which is quite costly since processes such as model training cost significant computational time. This limits the efficiency of node exploration. To address these challenges, we introduce a hybrid reward mechanism for node evaluation which combines LLM-estimated evaluations with actual performance scores to guide the search process more efficiently. For each node $x$, we employ an LLM-based value model $M_{\text{value}}$ to predict its potential performance. The value model takes as input the solution $\sigma_{\text{sol}}(x)$ and outputs an estimated score $s_{\text{LLM}}(x)$. This score is based on a comprehensive set of machine learning evaluation criteria, which is detailed in Appendix D.2. The LLM-estimated evaluation allows us to assess the quality of a node’s solution before conducting computationally expensive rollouts. This early evaluation helps prioritize nodes with higher potential, improving the efficiency of the search process. After the "simulation" of a node $x$, we obtain the actual performance score $s_{\text{actual}}(x)$ based on the development set, which reflects the true performance of the node’s solution. To smoothly transition from LLM-estimated values to actual performance scores, we implement an adaptive reward blending strategy. The Q-value $Q(x)$ for a node $x$ is updated as follows:
@@ 3.4 Tree Search Process
Upon simulation completion, the Q value is propagated backwards through the tree structure. This process updates the value and visit count of each parent node, from the simulated node to the root. Consequently, nodes associated with superior solutions receive higher priority in subsequent iterations. This allows nodes representing more promising solutions to be prioritized in future rollouts. Similar to Chi et al. (2024), I-MCTS implements a state-saving mechanism that caches the stage code for each node, which allows it to reuse previously generated code when a new configuration shares components with existing ones. For the detailed algorithm flow, please refer to Appendix  1.
@@ 4 Experiment
4.1 Experimental Details
For fair comparison, we follow Chi et al. (2024) to benchmark our method on 20 machine learning datasets, including 13 classification tasks and 7 regression tasks from the AutoML Benchmark (AMLB)  Gijsbers et al. (2024) and Kaggle Competitions. Five baselines are included for comparison, including AutoGluon Tang et al. (2024), AutoSklearn Feurer et al. (2022), AIDE Schmidt et al. (2024), Data Interpreter Hong et al. (2024a), SELA Chi et al. (2024). Experiments are conducted based on the MetaGPT Hong et al. (2024b) framework [^2] https://github.com/geekan/MetaGPT. All datasets are split into training, validation, and test sets with a 6:2:2 ratio. For the LLM-based Agents (I-MCTS, SELA, Data Interpreter, AIDE, etc), we choose Qwen2.5-72B-Instruct as the foundation LLM and employ a consistent initial task prompt across all methods  [^3] Details can be refer to our code repository https://anonymous.4open.science/r/I-MCTS. This prompt encompasses the dataset name, target column, and evaluation metric. We set the temperature parameter to 0.5. $\alpha_{\text{explore}}$ is set to 2. AIDE conducts 10 iterations per execution, while I-MCTS, SELA perform 10 rollouts. For I-MCTS and SELA, we employ Data Interpreter as the experimenter. The value of $\gamma$ in HRM is set to 0.2 across all datasets. This value is to control the decay rate of the LLM estimated score. Larger $\gamma$ makes the influence of the LLM estimated to persist longer, while smaller $\gamma$ allows the actual performance results to dominate the hybrid reward score more quickly. $\alpha_{explore}$ is set to 2. This aims to balance exploration and exploitation in the method’s search strategy. Following Chi et al. (2024), each method, except for AutoGluon, is run three times for each dataset. AutoGluon, being deterministic, is run only once with its default settings. AutoSklearn is also run with default settings, limited to 600 seconds per task.
4.2 Evaluation Metrics
@@ 4.6 Case Study
We conduct a comprehensive visualization analysis of the tree search process using the kick dataset. As show in Appendix E, our empirical results reveal that I-MCTS effectively leverages the introspective information from previous nodes to generate task-specific and actionable insights, while SELA exhibit significant homogeneity and lack specificity. Furthermore, our hybrid reward mechanism enhances the exploration efficiency, enabling more effective identification of high-quality nodes within the same computational budget.
@@ 5 Conclusion
While the proposed I-MCTS approach demonstrates significant improvements in AutoML agent performance, several limitations remain that warrant further investigation. First, the computational overhead of the introspective process, although beneficial for enhancing decision-making, introduces additional resource requirements. This limits the scalability of the method, particularly in scenarios where computational resources are constrained. Future work should explore optimizations to reduce this overhead and investigate how the approach scales with increased computational power. Also, our current implementation limits introspection to parent and sibling nodes to strike a balance between reflective depth and computational feasibility. Exploring deeper introspective hierarchies or incorporating a longer-term memory Xu et al. (2025) of the search trajectory are fascinating and important next steps. Second, the current evaluation of I-MCTS is primarily focused on structured data and tabular ML tasks. Its effectiveness on more complex and heterogeneous data types, such as image and text data, remains unexplored. Extending the application of I-MCTS to these domains could provide valuable insights into its generalizability and robustness across diverse machine learning tasks, such as Vision, NLP and RL tasks.
@@ Appendix A Algorithm
0: Problem description $p$, dataset $D$, metadata $d$, LLM LLMLLM, rollout number $k$, blending constant $\gamma$.
@@
[…]
@@ Appendix B Detailed Results
Click_prediction_small 26.6 26.6 40.2 40.3 26.1 39.4 12.9 13.9 10.1 10.4 29.1 30.2
GesturePhaseSegmentationProcessed 69.3 69.3 67.2 68.4 56.3 68.1 60.1 64.4 63.0 63.4 67.2 68.5
Moneyball 24.3 24.3 13.1 13.8 23.8 24.6 9.5 24.5 41.2 41.6 41.6 41.7
SAT11-HAND-runtime-regression 12.6 12.6 10.3 10.3 12.0 12.1 11.4 11.9 23.3 23.7 21.7 24.6
boston 39.8 39.8 19.5 19.6 40.5 41.3 37.0 38.6 62.1 62.5 62.5 62.7
colleges 88.3 88.3 2.1 2.1 86.0 87.8 87.5 87.7 93.8 94.1 94.1 94.3
concrete-strength 28.3 28.3 17.4 17.9 28.3 28.3 28.8 29.6 46.8 47.3 47.3 48.2
credit-g 50.5 50.5 35.1 44.0 21.6 48.4 48.1 53.2 40.1 40.8 50.9 51.1
diamonds 13.8 13.8 8.7 8.7 13.7 13.7 13.5 13.6 26.4 26.7 26.8 26.8
house-prices 9.0 9.0 2.0 2.0 8.9 8.9 8.5 9.0 17.9 18.2 15.8 18.4
icr 76.2 76.2 70.4 79.2 31.7 35.9 57.8 60.6 71.8 72.3 72.3 79.0
jasmine 84.3 84.3 84.4 84.7 83.6 84.6 77.8 83.5 83.0 83.5 84.2 84.3
kc1 38.3 38.3 43.5 45.0 40.8 42.6 38.1 41.2 38.1 38.6 44.8 45.6
kick 39.6 39.6 41.8 42.1 14.9 38.6 2.8 4.2 33.7 34.1 36.2 42.4
mfeat-factors 96.7 96.7 97.1 97.5 94.4 94.5 93.0 96.0 95.0 95.2 94.4 95.4
segment 93.5 93.5 92.7 93.1 91.7 92.2 91.7 92.6 93.2 93.5 91.1 93.7
smoker-status 78.0 78.0 78.6 78.9 74.8 76.3 77.3 81.5 85.8 86.4 79.9 91.5
software-defects 51.5 51.5 61.1 61.7 49.7 49.8 54.5 57.3 53.2 53.7 60.5 60.7
titanic 78.9 78.9 76.2 78.9 81.2 83.7 76.0 78.5 68.0 68.6 76.1 78.5
wine-quality-white 65.4 65.4 60.7 61.4 62.9 65.1 61.2 61.6 62.2 62.6 63.3 65.5
Overall NS % $\uparrow$ 53.2 53.2 46.1 47.5 45.5 51.8 47.4 50.2 55.4 55.9 58.6 59.8
@@
Appendix C Datasets
@@
Appendix D Prompts
D.1 Introspection Node Expansion Prompt
@@
You are an AI assistant tasked with analyzing a machine learning solution and proposing new insights to IMPROVE its performance. Given the current plan, solution code, Evaluation Metrics and performance on development score, suggest innovative approaches to ENHANCE and IMPROVE the performance.
@@
Based on this information, please propose ONE new insight w.r.t. the following specific stage of the machine learning pipeline to IMPROVE the overall performance. Your insight should be specific, actionable, and have the potential to IMPROVE the model’s performance.
@@
"insight": "ONE insight that specific, actionable, and have the potential to IMPROVE the model’s performance."
D.2 Node Evaluation Prompt
@@
Appendix E Case Study
@@
E.1 Overview of I-MCTS’s search process
@@ 99 Simulated: False
E.2 Overview of baseline SELA’s search process